Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
sorrentino - 13:07 Tuesday 07 May 2024 (64199) Print this report
Adjusting the B1p camera central image

This morning between 8:00 and 10:25 UTC I slightly changed the position of the central crop in B1p camera image, in order to better catch the central spot. See previous crop in first attachment (in LN3), current one in second attachment (in arms locked)

Images attached to this report
Comments to this report:
sorrentino - 9:53 Wednesday 15 May 2024 (64263) Print this report

When SR mirror is misaligned the central spot in B1p image becomes much fainter than the remaining part of the beam. The three images in attachment are with SR aligned, mid way during a typical locking sequence, and with final misalignment respectively.

The first plot in attachment shows the power evolution on B1p photodiode and on the central crop in the B1p image during the same lock acquisition. The power trend in the central area is quite noisy and seems not to give useful information about TCS tuning, see last pot.

Images attached to this comment
Search Help